OpenStudy (anonymous):

Somebody please help me solve this, or tell me how to do the steps? BEST ANSWER RECEIVES A MEDAL AND I WILL BECOME A FAN! Solve for x: 2 + x + 10 = 13

OpenStudy (anonymous):

simplify the problem (add like terms): 2+10=12 x+12=12 subtract 12 from both sides: x+12=12 -12 -12 x=0

OpenStudy (anonymous):

Thankyou so much! <3